a store keeper has cashew nuts that sell for $130 per kilo gram and peanuts that sell for $60 per kilogram. he wishes to make a mixture of 50 kg of the two kinds of nuts so that he can sell it for $78 per kg. how many kgs of each kind of nuts does he need to put in the mixture?
Answer by mananth(16946)


cashew nuts that sell for $130 per kilo gram let quantity be x
peanuts that sell for $60 per kilogram ---------- (50-x)
Mixture 50 kg at $78 /- kg
130x +60(50-x) = 50*78
130x +3000 -60x = 3900
70x =900
cashew nuts 12.9 kg
peanuts = 37.1 kg